# A Meat-Summer Night’s Dream: A Tangible Design Fiction Exploration of Eating Biohybrid Flying Robots

### v1

@@ 3.6. Post-Event Interviews and Data Analysis
After the immersive dining event, semi-structured individual interviews were conducted with each participants. The interview questions aimed to gather participants’ reflections on their overall experience of the drama and dinner event, with a focus on both the sensory and ethical dimensions of the staged future. Participants were asked to share their thoughts on the drama itself and provide opinions on key topics, including philosophical, ethical, environmental, economic, and socio-cultural issues from PESTEL Framework (Yüksel, 2012), which is a framework or tool used by marketers to analyze and monitor the macro-environmental impacts on an organization, company, or industry (Yüksel, 2012). It examines the Political, Economic, Social, Technological, Environmental, and Legal factors in the external environment. Additionally, we explored their perceptions of the identity of the animal-robot hybrid and their views on AI-infused food systems.
A thematic analysis was conducted to identify and interpret key patterns (Braun and Clarke, 2012). First, the transcripts were reviewed to familiarize with the data, and initial codes were created based on recurring topics. These codes were then grouped into broader themes that were were used to develop and organize the result section below, responding to our overarching research questions.
Although the participant group was small, the goal of this study was not to generate generalizable claims, but to probe diverse reactions and reflections using a depth-over-breadth approach common in design research and speculative methods.
4. Results and Discussion
In this section we present results alongside discussion, situating findings within HCI, HRI, and food futures research. Participants expressed curiosity, caution, and concern about the role of biohybrid robotics in reimagining food. Our analysis identified three themes: Technological Feasibility and Pathways, Cultural and Ritual Transformation, and Ethics and Moral Boundaries. Each integrates multiple perspectives, capturing points of convergence and divergence. We also consider how participants engaged with the research method itself, reflecting on the sensory cross-media design and the productive role of serendipity.
These reflections are not predictions or prescriptive ethical claims. Rather, they represent situated perspectives that illuminate a contested terrain where hopes, anxieties, and imaginaries of possible food futures—biohybrid or otherwise—circulate and collide.
@@ 4.1.1. Technological Feasibility and Pathways
As non-experts, participants offered varied yet complementary perspectives on the maturity, timelines, and scalability of biohybrid food technologies. Several viewed lab-grown meat and partial integrations—such as meat grown on robotic substrates—as plausible in the near term, while fully integrated biohybrid robots producing edible tissue were imagined as a more distant possibility. P6 suggested that building an autonomous biohybrid animal with an AI “brain” would be prohibitively costly but not inconceivable, estimating a 40-year horizon before such systems might emerge. In the meanwhile, partial implementations will more likely emerge and serve as marketing tools hoping to drive early adoption. By contrast, P3 drew on science fiction imaginaries, envisioning similar futures within a decade but stressing that adoption should be slowed to allow society to grapple with social and ethical complexities. P4 identified a more immediate and sensorial roadblock—improving the taste of lab-grown meat—arguing that if taste improves, adoption would accelerate and the technology upscaled. They speculated that market pressures might drive innovation toward experimenting with novel taste and nutrient profiles or even with cross-fertilisations of ideas and materialities—e.g., bringing together different species, body parts, and biological processes into a new hybrid formation—such as producing “tiger eggs”. While speculative, these anticipations highlight deep uncertainties about the societal, ethical, and disruptive potential of biohybrid robotics, echoing debates within the field itself (Mestre et al., 2024).
Several participants also questioned whether synthetic food production is as environmentally sustainable as often claimed, particularly compared to plant-based alternatives. Concerns included the resource intensity of engineering food at scale (P1), the energy demands of biohybrid and AI-driven processes (P2, P5), and the broader ecological footprint relative to traditional farming (P5). P4 expanded the debate by suggesting sustainability might also be measured in terms of disease reduction, such as lowering zoonotic risks from wild animal consumption. These exchanges underscore the difficulty of defining “sustainability” for emerging food paradigms, especially when both technologies and practices remain unsettled. Crucially, these tensions also speak to a wide set of risks that must be navigated moving forward. Importantly many of the associated risks (i.e., human health, AI domination, and environmental risks) have been key features of bioethical inquiry in tangent and more established fields—such as synthetic biology (see (Kirksey, 2021)).
Market adoption was another recurrent theme. P4 framed diffusion through psychographic segmentation and technical customization. Initially imagining biohybrid products as “a blob…a hamburger patty” he revised expectations toward modularity and tuning, “You might have a bunch of these [biohybrid robots] and they might be customized…affect the flavor or umami…with the nutrients that go into the robot-grown part…so it would make sense to use artificial intelligence as part of the formula.” He envisioned a staged diffusion despite high costs, “Early adopters will…try it anyway…then the fast follower markets like experiencers who would do it for the thrill…whereas the more information-seeking market…will be more interested in sustainability, environmental impacts…[and] whether it fits into their lifestyle.” He also added a status-oriented tier (“achievers”) who “watch the other markets” and adopt once it “makes them look good.” These psychographic distinctions, he argued, would vary across cultural contexts.
Beyond consumer markets, participants speculated on system-level effects. P4 pointed to food supply chains and the political stakes for governments choosing between promoting innovation or protecting traditional farming. P6 extended customization to the individuality of biohybrid animals, envisioning ranchers as curators who fine-tune traits such as personality and taste for differentiated offerings. P5 imagined personalization at the consumer end, where families could assemble or breed unique variations, creating heirloom-like recipes passed down across generations and producing affective attachments that transcend efficiency or sustainability.
Scaling pathways were also explored. P2 invoked a “Tesla model” in which early, costly applications give way to affordability as supply chains mature, “There were early adopters with very specialized expensive tasks, and then gradually a supply chain was built up to make it more affordable to produce.” He also referenced MAYA (Most Advanced Yet Acceptable) to stress that designs balancing familiarity and novelty can smooth transitions from niche enthusiasm to mainstream uptake. While Rogers’ diffusion framework emphasizes categories like innovators and laggards (Rogers, 2003), our findings show how lay participants developed imaginaries from such categories to articulate culturally grounded adoption models, informed as much by sensory and ethical reflection as by perceptions of technical readiness.
These reflections suggest that speculative dining can surface adoption dynamics often overlooked by surveys or foresight methods. Prior HFI work has highlighted how food technologies augment sensory and social experiences (Khot and Mueller, 2019; Obrist et al., 2014b; Deng et al., 2023b), but our study extends this by showing how embodied encounters provoke deeper reasoning about systemic issues—supply chains, market drivers, scaling transitions and regulation. This resonates with speculative design research (Dunne and Raby, 2024; Auger, 2013), which emphasizes provocation, yet here participants did more than imagine possibilities: they actively modeled barriers and accelerators of adoption. They did so with an acute sensitivity to a range of cultural and political themes, that we turn to now.
4.1.2. Cultural and Ritual Transformation
Participants imagined divergent futures for biohybrid food technologies, reflecting both optimism and deep concern. P4 sketched a dystopian scenario in which only the wealthy gain access, while climate change and resource scarcity exacerbate inequities, forcing migration and conflict. In contrast, their optimistic vision imagined widespread access, efficient production, and support for small farms and local communities. Such visions exemplify the splintering of sociotechnical imaginaries “towards utopic and dystopic extremes” (Barker and Jewitt, 2022), highlighting that possible futures for these technologies already operate within contested political and cultural spaces. Below we trace these opposing politics through the forces and metaphors outlined in Table 2.
Table 2. Conservative vs. Progressive Politics
Conservative Politics Progressive Politics
Lag – Cultural, political, or ethical frameworks trail behind technological developments, slowing acceptance and integration. Spark – Technology or cultural shifts ignite enthusiasm to embrace and accelerate adoption.
Drag – Institutions or communities that resists or slows innovation, creating structural friction or inertia. Rupture – Political, cultural and artistic movements that anticipate future transformations, preparing the ground for change.
Preservation – Efforts to use new technologies to maintain existing traditions, values, or rituals. Adapting only superficially. Transformation – Making new social, ethical, and cultural practices with emerging technologies and possibilities that depart from the old.
Even if the field of biohybrid robotics rapidly accelerates, participants emphasized that cultural acceptance may lag behind technological development. This tension became especially apparent when participants explored how future populations might respond to autonomous animal-like behaviors and forms in synthetic food—features intentionally introduced through the Flying Drone Performance. Discussions around creating “lifelike” robotic food from quasi-living materials often turned to the influence of political, religious, and cultural dynamics on adoption or rejection. For example, P5 explained: “Beef is such a politically charged topic in India because cows are sacred to Hindus.” She speculated that lab-grown beef could, in principle, enable those who wish to eat beef to do so while respecting religious doctrine. Yet she doubted whether technological innovation alone could bypass such deeply rooted sensitivities, asking “If [the cow is] lab-grown… does it solve all our problems?” This rhetorical question signals a hesitation where ingrained cultural attitudes are perceived as anchors that drag, slowing and resisting the adoption of biohybrid food technologies.
Other participants suggested biohybrids might be mobilized to preserve rituals and traditions. The research design itself staged this through the dinner-in-the-drama, where a flying robot evoked the banned French culinary practice of eating Ortolan Bunting—a practice now legally and ethically untenable. P2 extended this preservation logic by noting that religious and ceremonial diets have historically adapted to new conditions while retaining symbolic resonance.
In contrast, some framed biohybrids as carrying progressive inertia. Understood as part of a “food revolution” (Califf and Mayne, 2022), they were positioned as capable of reshaping culinary practices, rituals, and aesthetics. Participants described two progressive dynamics: gradual cultural shifts and sparks of transformation.
P1 emphasized familiarity as a bridge. She argued that biohybrids would initially join existing rituals, adopting the “trappings of traditional food” before evolving into distinct forms: “Someone invented ‘the hamburger’… Someone will invent the hamburger of synthetic meats [and biohybrid robots]. It’s going to be different.” Drawing parallels to early design strategies—trolley cars made to resemble trains, or push-button elevators with attendants—P1 suggested that cultural adjustment would first be eased through mimicry, before giving way to new aesthetics, flavors, and rituals.
Others stressed rupture and spark. P6 speculated that individuality in biohybrid animals could become a site of cultural experimentation, where advertised “personalities” provoke new attachments and ethical dilemmas. Conversely, if mass-produced and indistinguishable, biohybrids might “feel more like robots”, evoking less moral concern. Such reflections illustrate how the characteristics of the technology itself can catalyze ethical and symbolic negotiation. Speculative dining thus served as a stage not only for debating sustainability but also for surfacing affective and political economies of eating. Participants’ reflections resonate with broader debates in food ethics and cultural studies (Appadurai, 1981; Probyn, 2016). They also point to HCI challenges: designing technologies that help publics navigate the trade-offs of both conservative and progressive forces. Cultural experimentation around synthetic meat is already visible in art and design—such as Catts and Zurr’s Disembodied Cuisine and Semi-Living Steak projects (Catts and Zurr, [n. d.]), which used semi-living tissue to provoke visceral responses. Participant engagement in this study likewise suggests that speculative food artifacts prompts critical explorations that can help craft new sociotechnical imaginaries. Deployed iteratively, through co-design processes, it is possible that novel technologies and practices could be developed, embraced, or resisted.
Across these accounts, participants revealed how biohybrid food technologies occupy an ambivalent cultural and political terrain, oscillating between drag and spark, preservation and transformation. From an HCI perspective, these findings extend Human–Food Interaction (HFI) research that has emphasized edible interfaces, playful augmentations, and choreographic food interactions (Tanaka et al., 2011; Murer et al., 2013; Khot and Mueller, 2019; Obrist et al., 2016; Deng et al., 2023b). Much less attention, however, has been given to the creative future of designing “with food” (see 2.1.1). As this section and Bourdieu (1984) reminds us, taste functions as political and cultural distinction. Our study shows how speculative food artifacts—such as biohybrid drones—can provoke reflection not only on how food tastes, but on what taste means culturally, while exploring alternative and speculative drone functions as food that extend the notions of possible human–drone interaction (Wang et al., 2024). Indeed edible robots (RoboFood, nd) and “computational” foods (Deng et al., 2024) may be set to remake the palate, introducing new flavors and textures. However doing so will inevitably reshape categories of rarity, luxury, and cultural/political distinction. Whether easing continuity or sparking rupture, biohybrids appear as catalysts for reimagining the rituals, aesthetics, and politics of food. These insights prepare the ground for the next section, which turns to participants’ ethical deliberations on life, death, and consumption.
4.1.3. Ethics
@@
Ethical considerations emerged as a central axis of reflection, with participants repeatedly returning to three interconnected debates: (i) sentience and suffering, (ii) individuality and symbolism, and (iii) replicability and losing control. They posed ethical tests, recalled personal anecdotes, and drew on imaginaries from adjacent fields such as synthetic biology, robotics, and AI. Across these debates, participants wrestled with how biohybrids may provoke new dilemmas as they continue the transition from “science fiction into real science and engineering” (Webster-Wood et al., 2022). These discussions situated edible (biohybrid) robotics within wider debates about technological agency, taboos, and moral boundaries.
@@
Sentience and Suffering. P6 observed that humans routinely anthropomorphise machines and animals, projecting sentience and emotion onto them—a phenomenon well documented in robotics (Duffy, 2003) and animal studies (Morris, 1986). For him, this raised a central objection: “We would be building a sentient being to kill it… there would be more public resistance.” P1 extended this, shifting responsibility to designers, “If we are making these things more similar to living organisms… at what point do they just approach the same organisms we are trying to replace and prevent suffering?”
@@
The group’s focus mirrors ongoing controversies in synthetic biology, where frameworks such as the 14-day embryo rule hinge conceptually on the onset of nociception and neurological development (Hyun et al., 2016; for Stem Cell Research, 2021). P3 broadened the debate by invoking religious worldviews in which all living things possess a soul. Prompted by the robotic bird’s unpredictable movements, she extended this to AI-driven or biohybrid systems: “Could they embody a form of soul?” P2 noted that resolving such questions requires “a philosophical position on the soul and materialism” to judge whether biohybrids might possess sentience and thus the capacity to suffer.
@@
Perhaps recognizing the intractability of such ontological debates, P2 offered an ethical test: if a biohybrid “looks like an animal” but “does not have feelings” harming it would be “ethically neutral”. He extended this provocatively, “Is it unethical to grow a human and eat the human? It’s not a feeling person… there’s no ethical harm.” While framed hypothetically, participants reacted with visible discomfort. P2 himself acknowledged the idea as “powerful and gross” illustrating how disgust functions as a moral boundary. These responses echo Kass’s notion of the “wisdom of repugnance” (Kass, 1997) developed through a bioethical investigation of (human) cloning: visceral reactions that mark limits where rational justification falters.
Individuality and Symbolism. Participants also reflected on how individuality might reshape the ethics of consumption. P6 speculated that if biohybrids displayed distinct personalities, acceptance and regulation would be far more complex: “If all of these biohybrids had their own… personalities, then people seem to be more wary politically.” P5 added that even subtle cues shift moral perception: “When the tissue is dependent on a scaffold, it feels more like a robot… but when it flies, it feels more like a bird, as if it will have free will.”
Here, debates on sentience became entangled with symbolism. P5 argued that meaning often outweighs materiality: “Sometimes it doesn’t matter whether it’s real or synthetic… it’s just the fact that this is a symbol of my faith or my religion or whatever, like my beliefs.” P6 illustrated this through a personal anecdote: his grandmother once killed a cat that ate rare cheese in rural Alaska—an act framed not by abstract principle but by survival. These exchanges highlight the need for relativist ethics that attend to context, symbolism, and cultural values. For future systems to be perceived as ethical, designers must account for how perceptions of individuality and symbolism shape moral status.
@@
Replicability and Losing Control. P1 warned that creating self-replicating biohybrids could introduce runaway effects. While some Xenobots have already demonstrated replication (Gómez-Márquez, 2023), achieving this in robotics would require AI-led integration of synthetic and biological structures. For P1, the risks were profound: “Life is a machine for self-perpetuation, so the odds of unintended consequences are so much higher for creating a living agent.” If capable of replication, biohybrids would enter “the category of any living being with an intrinsic drive to reproduce”, potentially acting like a virus against other systems. These anxieties connected to wider fears about AI, where distributed agency and emergent behavior defy governance. As P1 put it, “It connects to AI… fears about general artificial intelligence having runaway impacts.” Such reflections were framed as calls for humility, guardrails, and continuous evaluation. P5 echoed this, questioning “How much control humans should exert: should biohybrids have feelings, agency, or senses? How much autonomy should they be allowed to develop?”
The questions and reflections foregrounded in this section show how technological agency operates across ontological (sentience, free will), perceptual (individuality, symbolism), and sociological (replication, runaway risks) levels. Participants attributed surprising degrees of agency and semi-sentience to the drones. Some described malfunctions as drones “having a mind of their own”; others referred to an “AI-brain” or “soul of the machine”. Such statements challenge the anthropocentric orientation of much AI and robot ethics, which typically centers on humanoid robots or conversational agents. Instead, participants engaged with animal–machine hybrids as ethically significant others. This aligns with posthumanist theory (Braidotti, 2013; Wolfe, 2010) and animal studies (Haraway, 2008), which decenter the human as the sole locus of moral value and open space for considering responsibility, care, and cohabitation with more-than-human actors. Further, these perceptions resonate with work on machine consciousness (Scott et al., 2023; Zhang et al., 2023; Xu and Sar, 2018), suggesting that moral reasoning emerges when technologies are seen as partially autonomous. From a Science and Technology Studies perspective (Latour, 2005), these responses illustrate how publics are already negotiating moral boundaries in spaces where “animal” and “machine” overlap.

### v2

@@ 3.6. Post-Event Interviews and Data Analysis
After the immersive dining event, semi-structured individual interviews were conducted with each participant. The interview questions aimed to gather participants’ reflections on their overall experience of the drama and dinner event, with a focus on both the sensory and ethical dimensions of the staged future. Participants were asked to share their thoughts on the drama itself and provide opinions on key topics, including philosophical, ethical, environmental, economic, and socio-cultural issues from PESTEL Framework (Yüksel, 2012). Additionally, we explored their perceptions of the identity of the animal-robot hybrid and their views on AI-infused food systems. Interview questions are listed in Appendix A.
Interviews were audio-recorded and transcribed. Given the study’s small and exploratory sample, we employed an inductive thematic analysis approach (Braun and Clarke, 2012) to surface rich, situated insights. Two researchers independently familiarized themselves with the transcripts and generated initial codes, ensuring analytical depth and reducing individual bias. The researchers then met to compare, refine, and consolidate codes into preliminary themes, resolving discrepancies through discussion and returning to the data as needed. Emerging themes were iteratively cross-checked against field notes and observational data from the event to strengthen interpretive validity and maintain alignment with participants’ embodied experiences. This reflexive, multi-stage process supported a trustworthy and transparent interpretation of the qualitative material. The goal of this study was not to generate generalizable claims, but to probe diverse reactions and reflections using a depth-over-breadth approach common in design research and speculative methods.
4. Results
Overall, participants expressed curiosity, caution, and concern about the role of biohybrid robotics in reimagining food. In our critical exploration of alternative food futures (see 4.1), we identified three themes: Technological Feasibility and Pathways; Ethics and Moral Boundaries; and Symbolism, Culture, and Rituals. Each theme integrates multiple perspectives, capturing both convergence and divergence. We also examine how participants engaged with the research method itself (see 4.2) through three lenses: Immersive Engagement and Emotive Responses; Multisensory Cross-Media Theatrical Design; and The Serendipity of Unexpectedness and Improvisation.
The reflections presented below are not offered as authoritative predictions or ethically prescriptive positions. Nor do they represent a unified vision for the aspirations surrounding emerging food technologies. Rather, they represent situated perspectives that illuminate a contested terrain where hopes, anxieties, and imaginaries of possible food futures—biohybrid or otherwise—circulate and collide.
@@ 4.1.1. Technological Feasibility and Pathways
Participants offered varied yet complementary perspectives on the maturity, timelines, and scalability of biohybrid food technologies. Several viewed lab-grown meat and partial integrations—such as meat grown on robotic substrates—as plausible in the near term, while fully integrated biohybrid robots producing edible tissue were imagined as a more distant possibility. Additionally, while some envisioned synthetic food technologies catering to niche markets, others anticipated broader economic implications, including challenges related to affordability and market accessibility.
P6 suggested that building an autonomous biohybrid animal with an AI “brain” would be prohibitively costly but not inconceivable, estimating a 40-year horizon before such systems might emerge. In the meantime, partial implementations are likely to emerge, serving as marketing tools to drive early adoption. By contrast, P3 drew on science fiction imaginaries, envisioning similar futures within a decade but stressing that adoption should be slowed to allow society to grapple with social and ethical complexities.
P4 highlighted taste as an immediate and sensorial barrier to adoption, arguing that improvements in flavor would accelerate acceptance and enable technological scaling. He suggested that market pressures could drive innovation toward technical customization—moving beyond familiar forms toward modular systems that allow tuning of flavor and nutritional profiles. This could involve experimenting with novel material combinations and biological processes, including cross-species or hybrid formations (e.g., “tiger eggs”), potentially supported by AI to optimize taste, umami, and nutrient composition.
P4 also envisioned a staged diffusion despite high costs, “Early adopters will…try it anyway…then the fast follower markets like experiencers who would do it for the thrill…whereas the more information-seeking market…will be more interested in sustainability, environmental impacts…[and] whether it fits into their lifestyle.” Complementing this view, P2 invoked a “Tesla model” of innovation, in which early, expensive applications enable the gradual maturation of supply chains and cost reductions.
Beyond consumer markets, participants speculated on system-level effects. P4 pointed to food supply chains and the political stakes for governments choosing between promoting innovation or protecting traditional farming. Regarding edible biohybrid robots, participants envisioned a dichotomy between homogenized mass production and artisanal, bespoke biohybrids with individual identities. This distinction structures divergent ethical and cultural consequences, which are examined in the following subsections and synthesized in Table 2. Extending this vision, P6 imagined customization at the production level, with ranchers acting as curators who fine-tune traits such as personality and taste to create differentiated offerings.
Participants also questioned whether synthetic food production is as environmentally sustainable as often claimed, particularly when considering its ecological impacts relative to other alternatives. Concerns included the resource intensity of engineering food at scale (P1), the energy demands of biohybrid and AI-driven processes (P2, P5), and the broader ecological footprint relative to traditional farming (P2). For instance, P2 criticized the current efficiency of lab-grown meat, stating that “It takes tons and tons of energy to do badly something that you’d use much less energy just raising a cow.” P5 extended the discussion by suggesting that understanding our relationship with animal–robot hybrids may be more important than their potential environmental impact.
Table 2. Individuality vs. Homogeneity in Biohybrid Food Futures
Dimension Individuality Homogeneity
Core Form Differentiated, customized, small-scale. Standardized, indistinguishable, mass-produced.
Market Analogy Small farms or locally sourced meat. Factory farming and industrial meat production.
Market Position Premium, bespoke products. Affordable, widely accessible products.
Governance Contentious, politically sensitive regulation. Compatibility with existing regulatory frameworks.
Human Perception Heightened lifelikeness; blurred animal–machine boundaries. Machine-like; reinforced technological status.
Ethical Implications Heightened ethical tension: perceived personality elicits empathy, complicating harm or consumption. Reduced ethical tension: commodification eases justification and regulation.
Food Culture Ritualized, sensory, affective engagement. Utilitarian, routine consumption.
4.1.2. Ethics and Moral Boundaries
@@
Ethical considerations emerged as a central axis of reflection, ranging from debates about sentience and relativist ethics to extreme speculative scenarios. Participants posed ethical tests and drew on imaginaries from adjacent fields such as synthetic biology, robotics, and AI. Across these discussions, they wrestled with how biohybrids might provoke new moral dilemmas, situating edible biohybrid robots within broader debates about technological agency, taboos, and moral boundaries.
@@
Sentience and Suffering. P6 observed that humans routinely anthropomorphize machines and animals, projecting sentience and emotion onto them. For him, this raised a central objection: “We would be building a sentient being to kill it… there would be more public resistance.” P1 extended this:
“If we are making these things more similar to living organisms… at what point do they just approach the same organisms we are trying to replace and prevent suffering?”
@@
P3 broadened the debate by invoking religious worldviews in which all living things possess a soul; she extended this to AI-driven or biohybrid systems: “Could they embody a form of soul?” P2 noted that resolving such questions requires “a philosophical position on the soul and materialism” to judge whether biohybrids might possess sentience and thus the capacity to suffer.
@@
Acknowledging the intractability of such ontological debates, P2 offered an ethical test: if a biohybrid “looks like an animal” but “does not have feelings”, harming it would be “ethically neutral”. He extended this provocatively, “Is it unethical to grow a human and eat the human? It’s not a feeling person… there’s no ethical harm.” While framed hypothetically, participants reacted with visible discomfort. P2 himself acknowledged the idea as “powerful and gross”, revealing the gap between harm-based ethical reasoning and deeply embedded social taboos that function as moral boundaries.
P2 further invoked Westworld and Ex Machina, noting how popular debates about machine ethics focus on humanoid robots while largely overlooking animal-like entities and biohybrids. In contrast, he emphasized that the “particular design [of the animal-robot hybrids] brought the conversation into the realm”.
Relativist Ethics and Tradeoffs. Participants reflected a relativist ethical stance in which the moral evaluation of biohybrid animals depends on context, form, and social framing rather than fixed principles. For example, P5 recounted a story of a grandmother in rural Alaska who killed and cooked a cat after it ate rare cheese, underscoring how scarcity and circumstance reshape moral boundaries: “It’s the context and the living environment where you are…for them it was survival.” This perspective extended to synthetic foods and biohybrids, where specific situations complicate simple distinctions between food and companion: “Do you think about it as food or do you think about it as an animal—why can’t it be both?”
P6 drew an analogy between polarized views on factory farming and locally sourced organic meat, suggesting that synthetic food could be framed along similar ideological lines, see Table 2. He expressed a preference for individualized biohybrid animals—akin to small-farm meat—arguing that their distinct traits and personalities heightened ethical sensitivity and made the experience more impactful: “If all of these animals had their own personalities, then people seem to be more wary.” P5 echoed that such individuality invites people to perceive biohybrids as subjects rather than interchangeable objects, fostering empathy, attachment, and a sense of moral responsibility that complicates purely instrumental or efficiency-driven reasoning around harm or consumption. By contrast, more homogeneous biohybrids were seen as eliciting less concern, as standardization and replaceability allow them to be framed as commodities or infrastructure, making large-scale production easier to justify.
P1 reflected on how ethical production practices might unintentionally diminish flavor, citing hothouse tomatoes as an example. She noted that tomatoes grown in stress-free greenhouse conditions often taste bland, whereas the sweetness and complexity of Mediterranean tomatoes emerge from environmental stressors such as uneven sunlight and pest exposure. This led her to question whether eliminating suffering—whether in plants or animals—might also erase some of the sensory richness people value, revealing a tension between ethical responsibility and culinary pleasure. Extending this idea, P6 described rare or unusual foods—such as bird’s nest soup, shark fin, or ortolan bunting—as “special” precisely because of their rarity and infrequent consumption, suggesting that novelty and scarcity can heighten sensory and emotional value even as ethical concerns remain salient.
@@
Replicability and Runaway Risks. P1 warned that creating self-replicating biohybrids could introduce runaway effects, considering the idea of AI-led integration of synthetic and biological structures in robotics. For P1, the risks were profound: “Life is a machine for self-perpetuation, so the odds of unintended consequences are so much higher for creating a living agent.” If capable of replication, biohybrids would enter “the category of any living being with an intrinsic drive to reproduce”, potentially acting like a virus against other systems.
These anxieties connected to wider fears about AI, where distributed agency and emergent behavior defy governance. As P1 put it, “It connects to AI… fears about general artificial intelligence having runaway impacts.” Such reflections were framed as calls for humility, guardrails, and continuous evaluation. P5 echoed this, questioning:
“How much control humans should exert: should biohybrids have feelings, agency, or senses? How much autonomy should they be allowed to develop?”
4.1.3. Symbolism, Culture, and Rituals
Participants noted that the symbolic dimensions of food technology shape both the preservation and transformation of culture and rituals. P1 emphasized familiarity as a bridge to help people adjust to innovation, expressing hope that food continues to carry cultural meaning rather than serving purely as sustenance. She argued that biohybrids would initially adopt the “trappings of traditional food” to ease novelty acceptance; over time, however, biohybrids would move beyond imitation, giving rise to new aesthetics, flavors, and rituals. Similarly, P2 referenced the principle of MAYA (Most Advanced Yet Acceptable) to stress that designs balancing familiarity and novelty can smooth transitions from niche enthusiasm to mainstream uptake.
Extending this discussion, P2 commented that the dinner theater revealed “a whole dimension about culture, cultural preservation, and cultural innovation”, describing the creation of the ortolan robot as “a positive and clever way to preserve a tricky piece of culture—using technology to ethically achieve what was once unethical or has since become problematic”. He further reflected on how cultural traditions can be preserved and ritual meaning maintained even as material forms change. He noted that religious and ceremonial diets have historically adapted to new conditions while retaining symbolic resonance, such as substituting animal sacrifice with ritual acts using bread. Through such adaptations, he argued, societies can sustain the meaningful dimensions of ritual while accommodating ethical, technological, or practical constraints.
In contrast, P5 noted, “Beef is such a politically charged topic in India because cows are sacred to Hindus.” She speculated that lab-grown beef could, in principle, enable those who wish to eat beef to do so while respecting those who oppose killing cows. Yet she doubted whether technological innovation alone could bypass such deeply rooted sensitivities, asking “If [the cow is] lab-grown… does it solve all our problems?” As she argued, cultural meaning often outweighs materiality:
“Sometimes it doesn’t matter whether it’s real or synthetic, it’s just the fact that this is a symbol of my faith or my beliefs.”
P5 also reflected on geographic diversity, pointing out that culinary innovation is often associated with specific locales—like Paris as a center of gastronomy, but noted that “other cultures with rich culinary history”, such as China or India, could equally serve as hubs for future food innovation, underscoring the interplay of local tradition, ethical concerns, and global imagination in shaping the future of eating.
Participants highlighted how biohybrid animals and synthetic meat can serve as media for cultural and artistic expression and as catalysts for cultural innovation. Distinct individuality in biohybrids was seen to enhance narrative engagement and provoke ethical reflection among diners. P6 suggested that by integrating the life stories and distinctive traits of biohybrids into dining experiences, restaurants could offer richer storytelling that engages both emotion and morality. Extending this idea to the domestic sphere, P5 envisioned personalization through family-led assembly or breeding of unique variants, producing heirloom-like recipes passed down across generations and fostering affective attachments.